# AdaptNC: Adaptive Nonconformity Scores for Uncertainty-Aware Autonomous Systems in Dynamic Environments

### v1

@@ 6.2 Social Navigation with Policy Shift
[…]
Table 1: This table contains the proposed evaluation metrics across all methods and case studies. AdaptNC achieves coverage closest to the target level of $90\%$ while maintaining low average uncertainty volumes. In contrast, competing methods exhibit a trade-off between coverage and volume, either attaining coverage targets with high volume, or missing them with small volumes. AdaptNC achieves the most stable coverage behavior over time, as evidenced by the lowest standard deviation of local coverage.
Performance Evaluation and Comparative Analysis: As reported in Table 1, AdaptNC achieves a global coverage of $90.00\%$, exactly matching the target coverage level of $90\%$, whereas all baseline methods exhibit undercoverage. Although Split CP produces the smallest uncertainty volume, it exhibits severe undercoverage, attaining only $5.98\%$ coverage and missing the target level by $84.02\%$. DtACI also fails to meet the target coverage, missing it by $16.5\%$, while incurring a significantly larger uncertainty volume than AdaptNC, amounting to $140\%$ of AdaptNC’s volume. This behavior reflects an inherent tradeoff between coverage and volume in DtACI arising from the fixed geometry of its nonconformity score. As shown in Fig. 4, baseline methods experience coverage degradation even before the collaboration radius increases at $t=2000$, highlighting the intrinsic difficulty of the multi-agent setting. Local coverage statistics in Fig. 3 further reveal intermittent coverage drops for the baselines, whereas AdaptNC maintains coverage consistently close to the target, remaining within one standard deviation of the local coverage target. In contrast, the ablation AdaptNC without Replay exhibits mild undercoverage while producing substantially larger uncertainty regions, corresponding to $160\%$ of AdaptNC’s volume. These results further validate the importance of the replay mechanism for achieving stable, efficient coverage.
6.3 Multirotor Tracking under Actuator Degradation
In this case study, we consider a multirotor tracking scenario in which a multirotor platform is tasked with tracking a figure-eight reference trajectory in the xxyy plane at constant altitude. Over the course of operation, the platform experiences stochastic actuator degradation, leading to a gradual loss of control effectiveness. The system is modeled using linearized $12$-dimensional multirotor dynamics following Dixit et al. (2023), where actuator degradation is captured through a multiplicative health matrix acting on the commanded control inputs. Each actuator’s effectiveness evolves stochastically, reflecting cumulative fault propagation commonly considered in predictive maintenance models (Shen et al., 2024). Trajectory prediction is performed using an LSTM trained under nominal conditions to predict future planar motion from a fixed window of past states. At test time, the growing mismatch between commanded and effective inputs induces a progressive shift in the system behavior, thereby violating the predictor’s implicit assumption of full actuation. Evaluation is conducted over a single continuous rollout of 6000 time steps under degradation, allowing the resulting prediction bias and policy shift to accumulate and providing a challenging benchmark for evaluating adaptive uncertainty quantification. Additional experimental details are provided in Appendix D.2.
Performance Evaluation and Comparative Analysis: As shown in Table 1, both AdaptNC ($90.17\%$) and DtACI ($89.65\%$) achieve coverage close to the $90\%$ target, however DtACI requires a region volume $920\%$ greater than that of AdaptNC. We see in Fig. 5 that shape adaptation enables significantly tighter conformal prediction regions while maintaining coverage. In contrast, Split CP exhibits pronounced undercoverage due to its non-adaptive calibration under distribution shift. The importance of replay is further illustrated by AdaptNC without replay, which shows systematic undercoverage and increased temporal variability in local and global coverage (Figs. 3 and 4). These results continue to highlight the importance of replay in stabilizing coverage under strong distribution shifts.
@@ 6.3 Multirotor Tracking under Actuator Degradation
Results Summary: Across all three case studies, AdaptNC maintains coverage close to the target level under structural distribution shifts while recovering significantly tighter uncertainty regions than competing methods. By jointly adapting the nonconformity score and threshold, AdaptNC avoids the conservatism of threshold-only adaptation and the undercoverage of non-adaptive baselines, with the replay mechanism mitigating coverage shocks and stabilizing local coverage. This behavior is also evident in the expert weight dynamics highlighted in Figures 1011, and 12, with AdaptNC exhibiting faster reweighting than DtACI and AdaptNC without Replay. Further details on this analysis could be found in Appendices D.2D.3, and, D.4.
@@ 7 Conclusion
We present AdaptNC, an online conformal prediction method that adapts nonconformity scores under distribution shift. AdaptNC leverages a reweighted data distribution to balance reliance on past data and introduces a replay mechanism to stabilize adaptation and maintain coverage as scores evolve. We establish theoretical regret and long-term coverage guarantees and validate the approach on three robotics case studies involving environmental change, multi-agent interaction, and system degradation. Across all settings, AdaptNC achieves coverage closer to the target level with lower uncertainty volume than competing methods, demonstrating its efficacy under distribution shift. Future work will address the limitations discussed in Appendix A.
Impact Statement
This paper presents a method for uncertainty quantification in autonomous systems operating under distribution shift. Our work is primarily methodological, focused on improving conformal prediction techniques, we recognize several potential societal implications. First, our method can improve the safety of the general public through enabling safer autonomy. The use of optimized non-conformity scores can help ensure that robots of the future offer high-performance through decreased conservatism while maintaining coverage.
@@ Appendix A Limitations
While AdaptNC demonstrates the ability to recover tight, valid uncertainty regions under distribution shift, it remains, as with other adaptive conformal prediction methods, susceptible to producing trivial uncertainty sets. Although such sets are not entirely uninformative, as they indicate that future observations are likely to fall outside the previously observed support and that the system may be undergoing a regime change, they nonetheless represent a limitation. To further examine this behavior, we report the percentage of vacuous coverage timesteps for each method in Table 2, defined as $\frac{1}{T}\sum_{t=0}^{T}\mathds{1}\{\hat{q}_{t,1-\alpha}=\infty\}$, which quantifies the frequency with which a method yields uninformative, unbounded uncertainty regions.
Method Indoor Localization Social Navigation Multirotor Navigation
AdaptNC 11.51% 40.42% 33.65%
AdaptNC w/o Replay 11.93% 24.52% 22.82%
DtACI 15.67% 21.17% 28.77%
Split CP 0.00% 0.00% 0.00%
Table 2: This table lists the percentage of timesteps exhibiting vacuous coverage for each method across all case studies. Methods incorporating temporal adaptation exhibit a nonzero frequency of vacuous coverage.
Table 2 shows that Split CP does not exhibit vacuous coverage at any time step, as it lacks the ability to adapt over time. In contrast, all methods incorporating temporal adaptation, including DtACI, exhibit a nonzero fraction of timesteps in which trivial uncertainty sets are recovered. Although AdaptNC achieves the lowest proportion of vacuous coverage timesteps for the Indoor Localization and Social Navigation tasks, the differences among the three adaptive methods are relatively modest. For the Multirotor Navigation task, AdaptNC attains the highest fraction of vacuous coverage timesteps; however, the remaining adaptive methods also display comparably high proportions. These results indicate that vacuous coverage is a shared limitation across all adaptive conformal prediction methods considered in this study.
Appendix B Method
B.1 Score Optimization Algorithm
@@
1: Generate high-density points $\hat{R}_{t}$ covering proportion $1-\alpha$ of density mass through Algorithm 4
@@ B.2 AdaptNC Algorithm
  Observe $X_{t}$, produce prediction $\hat{Y}_{t}$, observe $Y_{t}$
@@
   Compute distribution $\mathcal{H}_{t}^{\omega}$ using Eq. 4
@@
B.3 Dynamically Tuned Adaptive Conformal Inference (DtACI)
Algorithm 3 Dynamically tuned Adaptive Conformal Inference (DtACI) (Gibbs & Candès, 2024, Algorithm 2)

### v2

@@ 6 Experiments
ANC 90.6% $\pm$ 0.3% 27.76 $\pm$ 2.08 90.2% $\pm$ 0.3% 91.6% $\pm$ 1.2% 43.56 $\pm$ 6.00 91.4% $\pm$ 1.2% 92.5% $\pm$ 0.7% 0.04 $\pm$ 0.01 92.3% $\pm$ 0.7%
ANC NR 94.8% $\pm$ 0.8% 38.64 $\pm$ 3.67 94.4% $\pm$ 0.8% 87.9% $\pm$ 3.0% 73.47 $\pm$ 9.62 87.6% $\pm$ 3.0% 80.1% $\pm$ 2.5% 0.07 $\pm$ 0.02 80.0% $\pm$ 2.5%
DtACI 92.2% $\pm$ 1.8% 43.09 $\pm$ 4.46 91.8% $\pm$ 1.8% 80.0% $\pm$ 3.5% 52.67 $\pm$ 8.10 79.7% $\pm$ 3.5% 87.4% $\pm$ 1.7% 0.44 $\pm$ 0.01 87.2% $\pm$ 1.7%
Sp. CP 97.1% $\pm$ 0.9% 98.91 $\pm$ 3.12 96.7% $\pm$ 0.9% 3.0% $\pm$ 0.8% 0.71 $\pm$ 0.02 3.0% $\pm$ 0.8% 60.1% $\pm$ 1.6% 0.21 $\pm$ 0.00 59.9% $\pm$ 1.6%
Sh. CP 96.9% $\pm$ 1.1% 102.1 $\pm$ 7.0 96.5% $\pm$ 1.1% 3.0% $\pm$ 0.8% 0.70 $\pm$ 0.02 3.0% $\pm$ 0.8% 0.7% $\pm$ 0.2% 0.003 $\pm$ 0.0003 0.6% $\pm$ 0.2%
Abbreviations: ANC= AdaptNC; NR = No replay; Sh. CP = Shape CP; Sp. CP = Split CP; CP = Conformal Prediction.
6.1 Performance Evaluation and Comparative Analysis
Table 1 summarizes performance across all methods and environments. AdaptNC achieves coverage closest to the target level of $90\%$ while maintaining substantially smaller uncertainty volumes, indicating a favorable trade-off between coverage and efficiency. Furthermore, the low variance across 50 seeds shows that these trends remain consistent, demonstrating the robustness of the method.
Figure 2: This figure shows the evolution of local coverage over a 100-step sliding window for a representative seed. For readability, an exponential moving average is shown, with the raw data in a lighter color. AdaptNC exhibits the lowest variability in local coverage, indicating stable recovery of tight uncertainty regions under distribution shift.
Figure 3: This figure illustrates the evolution of empirical coverage over the evaluation horizon relative to the target coverage level (shown in red) for a representative seed. AdaptNC’s performance highlights that score-function adaptation enables sustained coverage with tighter regions.
Global Coverage and Local Coverage Stability: AdaptNC consistently achieves coverage closest to the target across all settings ($90.6\%$, $91.6\%$, and $92.5\%$) while maintaining stable local coverage with low variance. In contrast, baseline methods exhibit a clear trade-off between coverage, stability, and efficiency. DtACI attains near-target coverage in some settings, but only with significantly larger uncertainty regions. Conversely, methods that achieve small volumes, such as Split CP and Shape-CP in certain environments, suffer from severe undercoverage (e.g., $\approx 3.0\%$ in social navigation). Moreover, these baselines display pronounced temporal variations (Figures 2 and 3): AdaptNC without replay achieves reasonable global coverage only through oscillatory local behavior, while DtACI and Split CP exhibit intermittent coverage drops under distribution shift. These results indicate that baseline methods cannot simultaneously maintain accurate, stable coverage and efficient uncertainty regions, whereas AdaptNC achieves all three.
Uncertainty Volume: AdaptNC achieves smaller uncertainty regions while maintaining near-target coverage. For example, in multirotor tracking, AdaptNC attains comparable coverage to DtACI with an 1111x reduction in volume ($0.04$ vs. $0.44$). A similar trend is observed in indoor localization, where DtACI achieves reasonable coverage but at significantly higher volume ($43.09$ vs. $27.76$). In contrast, methods that achieve low volumes do so by sacrificing coverage, as seen with Split CP. Thus, baseline methods either achieve efficiency at the cost of coverage or achieve coverage with excessive volume. Figure 4 further illustrates that shape adaptation enables significantly tighter conformal prediction regions while maintaining coverage.
Importance of Replay: Figures 2 and 3 illustrate the evolution of global and local coverage over time for all case studies using a representative seed, highlighting the role of the replay mechanism. A consistent trend is observed across settings: AdaptNC maintains coverage closer to the target with significantly more stable behavior. In contrast, AdaptNC without replay exhibits increased variability in local coverage, which propagates to oscillations in global coverage and leads to more conservative uncertainty regions. These results highlight the importance of replay in stabilizing coverage under distribution shift and enabling the recovery of tighter, more accurate uncertainty regions.
Runtime Analysis: We report per-step runtimes in Table 2, averaged over 50 seeds on a desktop workstation (AMD Ryzen 7900X, 64GB RAM). Runtimes include all operations required to return a prediction region, including score evaluation (for offline methods Split-CP and Shape-CP). The threshold update in AdaptNC has a runtime comparable to DtACI, under 0.10.1ms in all environments. The score update averages $\approx 100$ ms, dominated by KDE evaluation, consistent with the theoretical analysis. Score updates are performed periodically (every $20$, $100$, and $100$ steps across the three case studies), keeping latency within the $t_{s}\cdot\Delta t$ budget. Importantly, periodic updates do not degrade coverage, as threshold updates occur at every timestep, which we verify empirically. Moreover, a runtime of $\approx 100$ ms is comparable to static safety filtering methods such as Control Barrier Functions (CBFs) [1], which must be recomputed in settings with distribution shift. Further discussion of these topics are provided in Appendices B.6 and B.7 respectively.
Table 2: Per-step runtime (ms) of all methods, averaged over 50 seeds.
Method Indoor Localization (ms) Multirotor (ms) Social Navigation (ms)
AdaptNC (threshold update) 0.09 $\pm$ 0.00 0.08 $\pm$ 0.00 0.08 $\pm$ 0.00
AdaptNC (score update) 100.61 $\pm$ 48.28 101.90 $\pm$ 38.06 97.34 $\pm$ 38.09
AdaptNC (no replay, threshold update) 0.15 $\pm$ 0.00 0.16 $\pm$ 0.01 0.13 $\pm$ 0.00
AdaptNC (no replay, score update) 7.04 $\pm$ 0.47 6.91 $\pm$ 0.12 6.55 $\pm$ 0.18
DtACI 0.09 $\pm$ 0.00 0.08 $\pm$ 0.00 0.08 $\pm$ 0.00
Split CP 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00
Shape CP 0.22 $\pm$ 0.06 0.20 $\pm$ 0.40 0.21 $\pm$ 0.04
Hyperparameter Sensitivity: We also provide a sensitivity analysis across different values of shape adaptation interval (update period) and replay buffer window size to assess the robustness of AdaptNC to moderate hyperparameter changes. Due to space constraints, the results are reported in Table 5 in Appendix E.5. Overall, AdaptNC exhibits low sensitivity to hyperparameters, with stable performance across a broad range of window sizes and update periods. Coverage and volume change noticeably only when deviating far from the base setting. The results show changes in the window size and update period have monotonic effects on both coverage and volume. The predictable nature of these tradeoffs enables straightforward, principled tuning.
Figure 4: This figure shows uncertainty regions from AdaptNC and DtACI at representative timesteps ($t\in[720,1560,3120]$) in the multirotor tracking task for a representative seed, along with the realized residual. AdaptNC recovers significantly tighter regions while maintaining coverage, whereas DtACI’s threshold-based adaptation constrains region geometry and leads to misalignment under distribution shift.
@@ 6.1 Performance Evaluation and Comparative Analysis
Results Summary: Across all three case studies, AdaptNC maintains near-target coverage under structural distribution shifts while producing significantly tighter uncertainty regions than competing methods. By jointly adapting the nonconformity score and threshold, it avoids the conservatism of threshold-only approaches and the undercoverage of non-adaptive baselines, with replay mitigating coverage shocks and stabilizing local coverage. This behavior is also reflected in the expert weight dynamics (Figures 1011, and 12), where AdaptNC exhibits faster reweighting than DtACI and AdaptNC without replay. Additional details are provided in Appendices E.2E.3, and E.4.
@@ 7 Conclusion
We present AdaptNC, an online conformal prediction method that adapts nonconformity scores under distribution shift. It leverages a reweighted data distribution and a replay mechanism to stabilize adaptation and maintain coverage as scores evolve. We establish regret and long-term coverage guarantees and validate the method on three robotics case studies involving environmental change, multi-agent interaction, and system degradation. Across all settings, AdaptNC achieves coverage closer to the target with lower uncertainty volume than competing methods. Future work will focus on improving computational efficiency and addressing the limitations discussed in Appendix A.
@@ Appendix A Limitations
While AdaptNC demonstrates the ability to recover tight, valid uncertainty regions under distribution shift, it remains, as with other adaptive conformal prediction methods, susceptible to producing trivial uncertainty sets. Although such sets are not entirely uninformative, as they indicate that future observations are likely to fall outside the previously observed support and that the system may be undergoing a regime change, they nonetheless represent a limitation. Empirically, AdaptNC exhibits a vacuous coverage rate of approximately $25\%$ across all case studies, while DtACI exhibits a rate of approximately $20\%$. Although AdaptNC shows a slightly higher vacuous coverage rate, the difference is marginal, and these results suggest that vacuous coverage is an inherent limitation shared across all adaptive conformal prediction methods considered in this study. Furthermore, we emphasize that vacuous coverage steps do not constitute a critical safety concern in autonomous systems. Rather, they serve as indicators of a potential regime change in the environment, at which point the system can transition to a fallback safe controller during the transient phase. We believe that score function regularization may be able to address this issue, and it is an avenue of future work we are pursuing.
Appendix B Additional Clarifications and Extended Discussion
B.1 AdaptNC is not a direct combination of DtACI and prior score-adaptation methods
While AdaptNC builds upon these important prior works, it is not merely a direct combination of these methods. The key novelty of AdaptNC lies in identifying and addressing a new setting that prior work does not consider: the joint, coupled adaptation of both the nonconformity score and the conformal threshold under distribution shift. This coupling introduces a non-trivial feedback loop. Changes in the score function alter the induced nonconformity distribution, which subsequently shifts the optimal quantile and destabilizes threshold updates. As a result, the update dynamics of prior independent methods no longer perform their function in the same manner. To support this new setting, we introduced an adaptive reweighting scheme that enables fully online score optimization, alongside new theoretical results including score function stability under joint adaptation (Theorem 5.3) in an asymptotic setting, using a guarantee on the stability of the score function adaptation in (Theorem B.1).
Furthermore, the direct amalgamation of DtACI with score adaptation, corresponds to our “AdaptNC without Replay” baseline, which includes our historical data reweighting scheme. As demonstrated in our ablation studies, this naive combination fails, exhibiting substantial misses in coverage and significantly inflated prediction regions because composition alone is insufficient to handle the feedback problem created through this joint optimization. The intuition behind this is presented in (Remark 5.5), which shows that there are two sources of shift in the score distribution; the bounded component due to the additional data point observed at the next timestep, and the potentially unbounded component which is due to the shift in the score function. To solve this, we introduced the counterfactual replay mechanism. By reinitializing and recalibrating the DtACI expert chain using recent data under the updated score, the replay mechanism prevents coverage shocks caused by mismatched quantiles. As shown empirically, replay is not a minor implementation detail but a fundamentally necessary component that restores both valid coverage and tight uncertainty sets.
Overall, AdaptNC goes significantly beyond combining existing techniques by formulating a new problem setting of coupled adaptation, identifying the instabilities of a naive synthesis, and introducing the novel mechanisms required to make joint adaptation both stable and effective.
B.2 AdaptNC’s Theoretical Support
AdaptNC’s various components have asymptotic guarantees. It is important to note that finite-sample guarantees are generally unavailable in online conformal prediction (CP) due to the absence of assumptions on distribution shift. Guarantees are often given in the asymptotic setting. [11] show that the regret of their expert weighting algorithm will have a bounded regret to the best expert. [12] establish asymptotic convergence of each expert to the target coverage. In this asymptotic stability setting, we guarantee that our method yields tight and consistent prediction regions. This is shown in Theorems˜5.2 and 5.3. Taken together, our method provides convergence guarantees in the asymptotic setting, including for the score adaptation procedure. Our theoretical analysis gives sufficient conditions under which replay induces a vanishing average perturbation of the ACI expert recursion. These conditions are asymptotic and rely on boundary anti-concentration; finite-sample replay guarantees remain an important direction for future work.
B.3 Role of the Low-Dimensional Assumption in Complexity Analysis
The assumption $d\leq 3$ in Section 5.4 is introduced to derive tractable bounds for the geometric operations underlying AdaptNC, particularly convex hull construction via the QuickHull algorithm. In low dimensions, QuickHull exhibits near-linear complexity, which supports the stated amortized per-timestep bound. This assumption is therefore analytical in nature and is used to characterize the regime where AdaptNC achieves favorable computational efficiency.
Importantly, AdaptNC itself does not depend on $d\leq 3$ and remains applicable in higher dimensions, although the associated geometric computations scale less favorably. In particular, QuickHull has complexity $O(n^{\lfloor d/2\rfloor})$ for $d>3$, where $\lfloor\cdot\rfloor$ denotes the greatest integer function. However, this growth remains manageable up to moderate dimensions (e.g., $d\leq 5$), where the scaling is at most quadratic, consistent with the empirical runtimes in Table 2. Moreover, in autonomous systems applications, which are the primary focus of this work, conformal prediction is typically performed over low-dimensional predictors such as scalar value functions or spatial trajectory predictors (up to $3$ dimensions), and thus AdaptNC remains applicable without incurring prohibitive computational overhead.
B.4 Effects of Vacuous Prediction Regions on the Next Cycle of Learning and Their Mitigation
The "infinite" safety margins or vacuous bounds do not corrupt the next cycle of learning, as the underlying data is not perturbed or destroyed. These vacuous bounds are often encountered when the system undergoes change, or sees significant shift in error patterns. When the non-conformity scores the system sees are very large, larger than those we have previously seen, AdaptNC correctly identifies that the system is undergoing change, and surfaces to the user that the change is major, and the system cannot yet accurately predict what will happen. In this way, vacuous bounds are informative; they show that AdaptNC cannot provide guarantees in high-change regimes. After this period of change is complete, AdaptNC will return to predicting safety regions, as the residuals stabilize. Finally, prediction of these vacuous bounds happens when the current quantile prediction $\bar{\alpha}_{t}$ becomes very small. The data is preserved in its unperturbed state in the history buffer $\mathcal{H}$. These independent mechanisms ensure that data is kept in a manner that permits learning from samples, even if the predicted safety margin is infinite.
B.5 Detecting Distribution Shift and Adapting to its Rate
Identifying when the data distribution has shifted is one of the major problems in the online conformal prediction setting. There are advantages to having more data, namely that more data allows for more data to develop bounds on the non-conformity scores. However, using old or stale data can yield inaccurate coverage which can lead to catastrophe for robots. AdaptNC is fundamentally a weighted mixture of a series of “expert" predictors. Each of these “experts" functions under the hypothesis that the distribution is shifting at a different rate, characterized by $\gamma$. A lower value of $\gamma$ indicates a low expected rate of change, and a higher $\gamma$ indicates a higher expected rate of change. AdaptNC uses the weights allocated to each expect to dynamically identify how quickly the distribution is changing, and use that estimate to change how much historical data it uses at every timestep. This is also illustrated in the expert weight dynamics (Figures 1011, and 12).
B.6 Ability to Maintain Safety if a Catastrophic Change Occurs Between Score Updates
AdaptNC separates adaptation into two components: The threshold update occurs at every timestep, while the score update occurs periodically. While the shapes update periodically, the threshold updates at every timestep. This provides instantaneous response to changes in the system, while the calculation of the new score function allows for AdaptNC to react to systemic changes in the system. We show empirically in our results that in scenarios with distribution shift, our method is able to maintain the target coverage better than the previous state of the art.
B.7 Tackling More Complex Tasks and Computational Considerations
While the task of drawing safety boundaries may appear computationally intensive, in practice it is not prohibitively slow. In our experimental settings, the score update path executed every $T$ steps takes approximately $100$ ms, while the threshold update path executed at every timestep takes approximately $0.1$ ms. The threshold update runtime is comparable to standard safety filtering methods such as Control Barrier Functions (CBFs), which typically assume negligible or no distribution shift. The score update enables adaptation to significant distribution shift. In such settings, recomputing CBFs is often more computationally demanding than the threshold update in AdaptNC.
Appendix C Method
C.1 Score Optimization Algorithm
@@
1: Generate high-density points $\hat{R}_{t}$ covering proportion $1-\alpha$ of density mass through Algorithm˜4
@@ C.2 AdaptNC Algorithm
  Observe $X_{t}$, produce prediction $\hat{Y}_{t}$
  Observe $Y_{t}$
  Append $(X_{t},Y_{t},s_{t})$ to $\mathcal{H}$
@@
   Compute distribution $\mathcal{H}_{t}^{\omega}$ using Eq.˜4
@@
C.3 Dynamically Tuned Adaptive Conformal Inference (DtACI)
Algorithm 3 Dynamically tuned Adaptive Conformal Inference (DtACI) [11, Algorithm 2]